[Halla_running] [New Logentry] RC Update Nov 30, 2021

xbai at jlab.org xbai at jlab.org
Tue Nov 30 18:25:01 EST 2021


Logentry Text:
--
<p><strong>Accelerator and Hall Status</strong></p><ul><li>Overall not much new, we are in stable production at present</li><li>Beam in Hall A is stable, we've been taking production runs with stable beam with few interruptions.</li><li>We did a ion chamber calibration in last swing shift, with SBS-ON, thin carbon target. This improved beam tripping situation.</li><li>Cycling through 2h LH2, 8h LD2 and 0.5 hour dummy target runs</li><li>Beam line optics will be kept as is.</li><li>A 2~3 hour long 10 uA and 12 uA on LD2 target run is planned in Tuesday Swing shift, this run includes the steps of gradually increasing beam current, see if GEM can tolerate it. RC will call GEM experts when this run happens</li></ul><p> </p><p><strong>Pending and New Issues:</strong></p><ul><li>Controlled access list for tomorrow:</li><li>BBCal and HCal: Arun, Sebastian, Ellen Becker with Scott as backup</li><li>GEM J0: Holly, Anu/Sean</li><li>Dave Kashy (work not related to SBS)</li><li>MCC!
 : on Moller PS</li></ul><p> </p><p><strong>Issues under process</strong></p><ul><li>Bob and David are working on a script to read scaler information, and calculate beam charge for each run, 90% finished</li><li>Eric is working on a simulation putting an absorber at the exit of the target chamber</li></ul><p> </p><p><strong>Subsystem Updates:</strong></p><ol><li>HCAL:Need to fix the raspberry pi controllers for HVs, and plug them into a controllable power supply; Updating Panguin plots</li><li>BBCal: Need to fix the trigger DAC GUI during controlled access</li><li>Timing hodoscope: None</li><li>Grinch: None</li><li>GEMS: A discussion on the risk of operating GEMs under 12 uA on LD2 target, we are doing it in today's swing shift, experts have been notified</li><li>Beamline: Code for accurate beam charge calculation is close to finish, 90%. looking for data to test.</li><li>DAQ: DAQ get stuck from time to time, probably due to MPD needs resetting regularly. Working o!
 n an alarm for DAQ when DAQ hungs</li><li>Replay analysis: Fou!
 nd an APV whose pedestal drifted from 400ADC to 900 ADC, probably need physical access to check if there's any hardware issue.</li></ol><p> </p><p><strong>Run Plan</strong></p><p>Follow the SBS-11 run plan listed on the wiki: <a class="external free" href="https://sbs.jlab.org/wiki/index.php/Main_Page" rel="nofollow">https://sbs.jlab.org/wiki/index.php/Main_Page</a></p><p> </p><p> </p>

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3953345
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halla_running/attachments/20211130/df605d89/attachment.html>


More information about the Halla_running mailing list